# Comparative Study of Multilingual Idioms and Similes in Large Language Models

### v1

@@ Abstract
Paria Khoshtab††thanks: Equal contribution, ordered alphabetically.1 Danial Namazifard11footnotemark: 11 Mostafa Masoudi1 Ali Akhgary1 Samin Mahdizadeh Sani1 Yadollah Yaghoobzadeh1, 2 1School of Electrical and Computer Engineering College of Engineering University of Tehran, Tehran, Iran 2Tehran Institute for Advanced Studies, Khatam University, Tehran, Iran {paria.khoshtab, namazifard, mostafa.masoudi, ali.akhgary, samin.mehdizadeh, yyaghoobzadeh}@ut.ac.ir
@@ 1 Introduction
Evaluating LLMs requires interacting with them, making prompt engineering a critical component for optimizing performance. We examine various prompting strategies, including chain-of-thought, few-shot, and dialogue simulation. We extend our evaluations to input being translated to English, as it serves as a strong baseline for many multilingual evaluations Lin et al. (2022); Shi et al. (2023); Liu et al. (2024b). To achieve a comprehensive evaluation, we conduct an exhaustive assessment using both closed-source—GPT-3.5, GPT-4o mini, and Gemini 1.5 Team et al. (2024)—and open-source Llama 3.1 (Dubey et al., 2024, 8B, 70B) and Qwen2 (Yang et al., 2024, 7B, 72B).
Our findings reveal several novel insights: (i) Prompt engineering methods show varying degrees of success depending on the figurative type, language, and model used. (ii) Open-source models perform similarly to closed-source models in most languages, but they struggle with low-resource languages in interpreting similes. (iii) The interpretation of idioms in the style of the MAPS dataset is nearing saturation for many languages when strong LLMs are used, due to the presence of idioms and their meanings in their training data. (iv) The presentation of pre-trained data, as well as the script used in different languages, significantly impact model performance (v) Chain-of-thought prompting proves particularly effective for simile interpretation in smaller models.
@@ 2.1 Figurative language processing
In addition to similes, proverbs and idioms are another type of figurative expression that has been explored in various studies, focusing on identifying whether a phrase is used idiomatically or proverbially, either within a specific context (token-level) or in general (type-level) Li and Sporleder (2009); Fazly et al. (2009); Verma and Vuppuluri (2015); Salton et al. (2016); Peng and Feldman (2016). Beyond detection and generation, researchers have examined methods for interpreting and representing these figurative expressions, including literal paraphrasing, treating them as single tokens, or composing them from characters rather than words Liu and Hwa (2016); Zhou et al. (2021). Additionally, significant research has been conducted on other types of figurative language, such as metaphors, sarcasm, and irony Yu and Wan (2019); Ghosh et al. (2017); Chakrabarty et al. (2021).
@@ 2.2 Multilingual prompt engineering
In addition to prompt engineering, translate-test methods are commonly employed to address multilingual challenges. In this approach, evaluation data is translated into English before processing, using tools like Google Translate or LLMs Ahuja et al. (2023); Liu et al. (2024a); Shi et al. (2023). This method has proven effective in reducing performance gaps across languages Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023). Building on this, Huang et al. (2023) propose Cross-Lingual-Thought (XLT) prompting, a technique that translates the question into English and solves the problem in English before generating a response in the original language. We study the impact of English translation on simile and idiom interpretation across several languages, evaluating the consistency of responses between original and translated inputs under zero-shot and CoT settings.
@@ 3.1 Datasets
We employ two existing datasets: MABL (Metaphors Across Borders and Languages) Kabra et al. (2023) for simile experiments, and MAPS (Multicultural Proverbs and Sayings) Liu et al. (2024b) to assess idioms. Both datasets facilitate the analysis of figurative language understanding across multiple languages, providing a diverse multilingual resource.
MABL contains figurative expressions in eight languages: English, Indonesian, Hindi, Swahili, Yoruba, Kannada, Sundanese, and Javanese. This dataset captures cultural and linguistic diversity in figurative language, offering a valuable resource for testing multilingual LLMs’ abilities. We randomly select 200 simile samples from each language for evaluation, as shown in Table 1.
@@
MAPS consists of proverbs and sayings, designed to evaluate their interpretation within conversational contexts. The dataset provides binary labels, indicating whether the proverb is used figuratively. It spans six languages: English, Indonesian, Mandarin Chinese, Bengali, German, and Russian, with sample counts of 214, 267, 143, 272, 183, and 226 respectively. We select idiomatic sentences from each language for evaluation, as detailed in Table 2.
@@
3.2 Persian datasets
In addition to using the existing datasets, we created two new datasets specifically for the Persian language, following the formats of MABL for similes and MAPS for idioms. These datasets provide resources for evaluating figurative language understanding in Persian, which is underrepresented in current research on multilingual models.
Simile
We follow the methodology used in the MABL dataset but utilize GPT-4o to assist in generating examples. The model generates simile pairs by producing a start phrase with two possible endings—one reflecting the correct simile interpretation and the other conveying an incorrect meaning. The Persian Simile dataset consists of 200 samples. The prompts closely follow the instructions from the MABL dataset, ensuring consistency with the format and objectives of the original dataset. After generation, three native Persian speakers manually evaluate and correct the examples to ensure both accuracy and cultural relevance.
Idiom
We follow a methodology inspired by the creation process of the original MAPS dataset. First, we collect Persian idioms from two online resources: Daneshchi 222daneshchi.ir, an educational portal, and Abadis 333abadis.ir, an online dictionary. To create conversational contexts based on each idiom’s explanation, we develop prompts using examples written by native Persian speakers to guide the model in understanding how idioms are used in everyday conversation. Using GPT-4o, we generate a conversational context for each idiom, ensuring that the idiom is correctly situated in a natural dialogue. In the second round, we provide the model with the idiom, its explanation, and the generated conversational context. The model is then tasked with generating two response choices—one correct and one incorrect—based on the meaning of the idiom in the conversation. The Persian Idiom dataset contains 316 samples. Finally, native speakers review the generated content to ensure its accuracy, cultural appropriateness, and grammatical correctness.
For more details on the dataset construction and verification process, see Appendix A.
3.3 Languages
@@ 3.4 Models
We evaluate the performance of several open-source and closed-source LLMs to understand their capabilities in processing figurative language across multiple languages. The models under consideration include GPT-3.5-Turbo-0125 OpenAI (2023), GPT-4o mini OpenAI et al. (2024), and Gemini 1.5 Flash Team et al. (2024) which are representative of closed-source commercial LLMs, as well as open-source models: Llama 3.1 (Dubey et al., 2024, 8B, 70B) and Qwen2 (Yang et al., 2024, 7B, 72B). These models are selected based on their widespread use and the contrast they offer in terms of accessibility, customization potential, and model size variations. Note that although Llama 3.1 does not cover all of our languages, it performed quite well in our initial experiments. We also use the default parameters such as temperature for all model configurations in all experiments for consistency.
We access these models through the APIs provided by OpenAI 444platform.openai.com, Gemini 555ai.google.dev, and OpenRouter 666openrouter.ai for accessing open-source models. The total cost for running the experiments is estimated to be under $40, with approximately $30 spent on OpenAI and $10 on OpenRouter.
@@ 3.5 Prompting
We use several promoting techniques with examples in native or English translation or combined. We explain the techniques and then mention which ones we used in native, English, and native-English setups. The instruction is consistently given in English across all settings, as shown in Table 3. The full prompt templates used in the experiments are available in Appendix B.
@@ 3.5.2 Input language
We follow the trend of translating either the training or test data into English (Shi et al., 2023; Conneau et al., 2018; Qin et al., 2023), utilizing three translation systems that cover all of our languages: the Google Translate API, Meta’s No Language Left Behind (NLLB-200-3.3B) model (Team et al., 2022), and GPT-3.5 OpenAI (2023). We investigate zero-shot and one-shot settings to explore the few-shot effect in this context. For CoT, we ask the model to explain the meanings of idioms or similes alongside a one-shot example in English to activate CoT reasoning abilities. For simile, we also experiment with dialogue simulation prompting techniques. For idiom tasks, as they are already presented in a dialogue format, dialogue simulation is not applied.
@@ 3.6 Evaluation Process
We utilize LLMs to evaluate both tasks. For each task, we select a specific LLM, apply the appropriate prompt, and retrieve the model’s answer based on its interpretation. The answers are parsed using regular expressions (regex) to extract the final binary result. In cases where regex fails to correctly extract the answer due to irregular formatting or model output variations, manual verification is employed to ensure accuracy in the evaluation process.
@@ 4.1 Native
On the other hand for closed-source models, CoT compared to zero-shot has improved efficiency for GPT-4o mini and GPT-3.5, though this trend is not consistently observed for Gemini 1.5. Additionally, CoT tends to be more effective for similes than idioms.
@@ 4.2 English translation
The results of using both native and English inputs (as explained in Section 3.5.2) are also shown in Figure 2 in the third column named “Both”. When the GPT-3.5 model is used, using both native and Translated-English is a more effective method in most languages in both figurative types. There are two exceptions in similes for Kannada and Persian, where the Translated-English approach performs better. This seems to be specific to certain languages and requires further investigation. When using Gemini 1.5, a more capable model, the results are similar to those obtained using the Translated-English method for both figurative types. We also observe that Gemini 1.5 never outperforms native prompting in idioms. This indicates that the model has a strong understanding of idioms in native, and translating them may lead to errors in interpretation.
@@ 5 Analysis
In this section, we present four further analyses of our observations to deepen our comparative study of simile and idiom interpretation.
@@ 5.1 The thought pathway in CoT
To further investigate the impact of language in CoT responses on simile interpretation, we conducted additional experiments. We explicitly prompted Gemini 1.5 to generate responses in the native language, and instructed GPT-4o mini to respond in English (Appendix D). While GPT-4o mini showed no significant performance differences when responding in English (except Sw), Gemini 1.5 displayed improved performance in Fa, Su, and Jv when responding in their respective native languages. Results from these new experiments are presented in Table 17 in the Appendix.
@@ 6 Conclusion
This study explored how multilingual LLMs interpret similes and idioms across various languages, comparing the performance of both open-source and closed-source models. By leveraging the MABL, MAPS, and our newly developed Persian datasets, we evaluated different prompt engineering strategies, including one-shot, chain-of-thought (CoT), and dialogue simulation. Our results demonstrate that while these techniques enhance performance, their effectiveness depends on the type of figurative language, the language in question, and the specific model used. Open-source models, such as Llama 3.1 and Qwen2, performed competitively overall but faced challenges in low-resource languages, particularly with simile comprehension. In contrast, idiom interpretation has reached near-saturation in most languages, indicating a need for more complex datasets to push the boundaries of LLM capabilities.
In addition, we introduced two new Persian datasets for simile and idiom interpretation, offering valuable resources for evaluating LLM performance in this language.
In conclusion, this research advances our understanding of figurative language processing in multilingual LLMs, pinpointing critical areas for improvement and future exploration. Moving forward, several key directions warrant further investigation. Expanding the scope of figurative language types to include metaphors, sarcasm, and irony could provide a more comprehensive evaluation of LLMs’ capabilities. Additionally, the creation of multilingual datasets that maintain consistency across languages would enable more robust cross-linguistic comparisons. Lastly, developing datasets that challenge LLMs with more context-dependent or ambiguous figurative expressions will be crucial for driving progress in this area.
@@ Acknowledgements
We extend our gratitude to Zheng-Xin Yong and Emmy Liu for their valuable feedback and insightful comments on an earlier version of this paper.
Appendix A Dataset Construction
@@ 1.1 Persian Simile Dataset
To create the Persian Simile dataset, we task GPT-4o with generating over 800 simile pairs, following the structure used in the MABL dataset. Each sample consists of a start phrase with two possible endings: one correctly reflecting the intended meaning and the other presenting an incorrect or opposite interpretation.

### v2

@@ Abstract
Paria Khoshtab††thanks: Equal contribution, ordered alphabetically.1 Danial Namazifard11footnotemark: 11 Mostafa Masoudi1, 2 Ali Akhgary1 Samin Mahdizadeh Sani1 Yadollah Yaghoobzadeh2, 1 1School of Electrical and Computer Engineering College of Engineering University of Tehran, Tehran, Iran 2Tehran Institute for Advanced Studies, Khatam University, Tehran, Iran {paria.khoshtab, namazifard, mostafa.masoudi, ali.akhgary, samin.mehdizadeh, y.yaghoobzadeh}@ut.ac.ir
@@ 1 Introduction
Evaluating LLMs requires interacting with them, making prompt engineering a critical component for optimizing performance. We examine various prompting strategies, including chain-of-thought, few-shot, and dialogue simulation. We extend our evaluations to input being translated to English, as it serves as a strong baseline for many multilingual evaluations Lin et al. (2022); Shi et al. (2023); Liu et al. (2024b). To achieve a comprehensive evaluation, we conduct an exhaustive assessment using both closed-source—GPT-3.5, GPT-4o mini, and Gemini 1.5 Team (2024a)—and open-source Llama 3.1 (Team, 2024b, 8B, 70B) and Qwen2 (Team, 2024d, 7B, 72B).
Our findings reveal several novel insights: (i) Prompt engineering methods show varying degrees of success depending on the figurative type, language, and model. (ii) Open-source models perform similarly to closed-source models in idioms, but they generally lag behind in interpreting similes. (iii) The interpretation of idioms in the style of the MAPS dataset is nearing saturation for many languages when strong LLMs are used, due to the presence of idioms and their meanings in training data. (iv) The presentation of pre-trained data, as well as the script used in different languages, significantly impact model performance. (v) Chain-of-thought prompting proves particularly effective for simile interpretation in smaller models.
@@ 2.1 Figurative language processing
In addition to similes, proverbs and idioms are another type of figurative expression that has been explored in various studies, focusing on identifying whether a phrase is used idiomatically or proverbially, either within a specific context (token-level) or in general (type-level) Li and Sporleder (2009); Fazly et al. (2009); Verma and Vuppuluri (2015); Salton et al. (2016); Peng and Feldman (2016). Beyond detection and generation, researchers have examined methods for interpreting and representing these figurative expressions, including literal paraphrasing, treating them as single tokens, or composing them from characters rather than words Liu and Hwa (2016); Zhou et al. (2021). In a concurrent study on idiom translation, Rezaeimanesh et al. (2024) introduce the PersianIdioms dataset—comprising parallel Persian-English and English-Persian samples, designed for evaluating idiom translation—and compare the performance of LLMs, NMT systems, and hybrid methods.
Additionally, significant research has been conducted on other types of figurative language, such as metaphors, sarcasm, and irony Yu and Wan (2019); Ghosh et al. (2017); Chakrabarty et al. (2021). Additionally, significant research has been conducted on other types of figurative language, such as metaphors, sarcasm, and irony Yu and Wan (2019); Ghosh et al. (2017); Chakrabarty et al. (2021).
@@ 2.2 Multilingual prompt engineering
In addition to prompt engineering, translate-test methods are commonly employed to address multilingual challenges. In this approach, evaluation data is translated into English before processing, using tools like Google Translate or LLMs Ahuja et al. (2023); Liu et al. (2024a); Shi et al. (2023). This method has proven effective in reducing performance gaps across languages Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023). Building on this, Huang et al. (2023) propose Cross-Lingual-Thought (XLT) prompting, that translates the question into English and solves the problem in English before generating a response in the original language. We study the impact of English translation on simile and idiom interpretation across several languages, evaluating the consistency of responses between original and translated inputs under zero-shot and CoT settings.
@@ 3.1 Datasets
We employ two existing datasets: MABL (Metaphors Across Borders and Languages) Kabra et al. (2023) for simile experiments, and MAPS (Multicultural Proverbs and Sayings) Liu et al. (2024b) to assess idioms. Both datasets facilitate the analysis of figurative language understanding across multiple languages, providing a diverse multilingual resource. A description of the datasets can be found in Appendix A; however, key points are explained in the following.
MABL
contains figurative expressions in eight languages: English, Indonesian, Hindi, Swahili, Yoruba, Kannada, Sundanese, and Javanese. This dataset captures cultural and linguistic diversity in figurative language, offering a valuable resource for testing multilingual LLMs’ abilities. We randomly select 200 simile samples (to be close to the number of examples in the idiom dataset, ensuring balance and comparability between the two datasets) from each language for evaluation, as shown in Table 1.
@@
MAPS
consists of proverbs and sayings, designed to evaluate their interpretation within conversational contexts. The dataset provides binary labels, indicating whether the proverb is used figuratively. It spans six languages: English, Indonesian, Mandarin Chinese, Bengali, German, and Russian, with sample counts of 214, 267, 143, 272, 183, and 226 respectively. We select idiomatic sentences for evaluation, as detailed in Table 2.
@@
Persian datasets
In addition to using the existing datasets, we created two new datasets specifically for the Persian language, following the formats of MABL and MAPS for similes and idioms that provide resources for evaluating figurative language understanding in Persian, which is underrepresented in current multilingual model researches.
(i) Persian simile We follow the methodology used in the MABL dataset but utilize GPT-4o to assist in generating examples. The model generates simile pairs by producing a start phrase with two possible endings—one reflecting the correct simile interpretation and the other conveying an incorrect meaning. The Persian Simile dataset consists of 200 samples. The prompts closely follow the instructions from the MABL dataset, ensuring consistency with the format and objectives of the original dataset (Appendix B). After generation, three native Persian speakers manually evaluate and correct the examples to ensure both accuracy and cultural relevance.
(ii) Persian idiom We follow a methodology inspired by the creation process of the original MAPS dataset. First, we collect Persian idioms from two online resources: Daneshchi 222https://www.daneshchi.ir, an educational portal, and Abadis 333https://abadis.ir, an online dictionary. To create conversational contexts based on each idiom’s explanation, we develop prompts using examples written by native Persian speakers to guide the model in understanding how idioms are used in everyday conversation. Using GPT-4o, we generate a conversational context for each idiom, ensuring that the idiom is correctly situated in a natural dialogue. In the second round, we provide the model with the idiom, explanation, and the generated conversational context. The model is then tasked with generating two response choices—one correct and one incorrect—based on the meaning of the idiom in the conversation. The Persian Idiom dataset contains 316 samples. Finally, native speakers review the generated content to ensure its accuracy, cultural appropriateness, and grammatical correctness.
For more details on the dataset construction and verification process, see Appendix B.
3.2 Language Categorization
@@ 3.3 Models
We evaluate the performance of several open-source and closed-source LLMs to understand their capabilities in processing figurative language across multiple languages. The models under consideration include GPT-3.5-Turbo-0125 Team (2023), GPT-4o mini Team (2024c), and Gemini 1.5 Flash Team (2024a) which are representative of closed-source commercial LLMs, as well as open-source models: Llama 3.1 (Team, 2024b, 8B, 70B) and Qwen2 (Team, 2024d, 7B, 72B). These models are selected based on their widespread use and the contrast they offer in terms of accessibility, customization potential, and model size variations. Note that although Llama 3.1 does not cover all of our languages (Just En, De, and Ru are covered), it performed quite well in our initial experiments. The cost for running experiments is given in Appendix C.
@@ 3.4 Prompting
We use several prompting techniques with examples in native or English translation or combined. We explain the techniques and then mention which ones we used in native, English, and native-English setups. The instruction is consistently given in English across all settings, as shown in Table 3. The full prompt templates used in the experiments are available in Appendix D.
@@ 3.4.2 Input language
We follow the trend of translating either the training or test data into English (Shi et al., 2023; Conneau et al., 2018; Qin et al., 2023), utilizing three translation systems that cover all of our languages: the Google Translate API, Meta’s No Language Left Behind (NLLB-200-3.3B) model (Team, 2022), and GPT-3.5 Team (2023). We investigate zero-shot and one-shot settings to explore the few-shot effect in this context. For CoT, we ask the model to explain the meanings of idioms or similes alongside a one-shot example in English to activate CoT reasoning abilities. For simile, we also experiment with dialogue simulation prompting techniques. For idiom tasks, as they are already presented in a dialogue format, dialogue simulation is not applied.
@@ 3.5 Evaluation Process
We utilize LLMs to evaluate simile and idiom tasks. For each task, we select an LLM, apply the appropriate prompt, and retrieve the model’s answer. The answers are parsed using regular expressions (regex) to extract the final binary result. In cases where regex fails to correctly extract the answer due to irregular formatting or model output variations, manual verification is employed to ensure accuracy in the evaluation process.
@@ 4.1 Native
On the other hand for closed-source models, CoT compared to zero-shot has improved efficiency for GPT-4o mini and GPT-3.5, though this trend is not consistently observed for Gemini 1.5. Additionally, CoT tends to be more effective for similes.
@@ 4.2 English translation
The results of using both native and English inputs (as explained in Section 3.4.2) are shown in Figure 2 in the third column named “Both”. When the GPT-3.5 model is used, using both native and Translated-English is a more effective method in most languages in both figurative types. There are two exceptions in similes for Kannada and Persian, where the Translated-English approach performs better. This seems to be specific to certain languages and requires further investigation. When using Gemini 1.5, a more capable model, the results are similar to those obtained using the Translated-English method for both figurative types. We also observe that Gemini 1.5 generally does not outperform native prompting in idioms. This indicates the model’s strong understanding of idioms in native, and translating them may lead to errors in interpretation.
@@ 5 Analysis
In this section, we present further analyses of our observations to deepen our comparative study of simile and idiom interpretation.
@@ 5.1 The thought pathway in CoT
To further investigate the impact of language in CoT responses on simile interpretation, we conducted additional experiments. We explicitly prompted Gemini 1.5 to generate responses in the native language, and instructed GPT-4o mini to respond in English (Appendix F). While GPT-4o mini showed no significant performance differences when responding in English (except Sw), Gemini 1.5 displayed improved performance in Fa, Su, and Jv when responding in their respective native languages. Results from these new experiments are presented in Table 17 in the Appendix.
@@ 6 Conclusion
This study examined how multilingual LLMs interpret similes and idioms across languages, comparing open-source and closed-source models using MABL, MAPS, and newly developed Persian datasets. We tested prompt engineering strategies like one-shot, CoT, and dialogue simulation, finding their effectiveness varied by figurative language type, language, and model. While open-source models like Llama 3.1 and Qwen2 performed well overall, they struggled with similes in low-resource languages. Idiom interpretation, however, showed near-saturation, highlighting the need for more challenging datasets.
Our two new Persian datasets contribute valuable resources for evaluating LLMs in this language. Expanding the scope of figurative language types to include metaphors, sarcasm, and irony could provide a more comprehensive evaluation of LLMs’ capabilities. Also, developing datasets that challenge LLMs with more context-dependent or ambiguous figurative expressions will be crucial for driving progress in this area.
@@ Acknowledgements
We extend our gratitude to Yong Zheng-Xin and Emmy Liu for their valuable feedback and insightful comments on an earlier version of this paper. We also sincerely thank the anonymous reviewers for their valuable feedback and thoughtful suggestions.
Appendix A Simile and Idiom Datasets
Here is a detailed explanation of the MABL and MAPS datasets, designed to evaluate figurative language understanding for similes and idioms, respectively.
1.1 MABL for Simile
The MABL dataset is designed to assess figurative language understanding, consists of similes. Similes are rhetorical devices that compare two different concepts, often using words like "as" or "like", to highlight similarities. Understanding similes requires a nuanced grasp of both literal and figurative meanings within a sentence, making this a challenging task for language models.
The primary task in the MABL dataset is a binary classification problem. Each instance, as shown in Table 1, consists of: start phrase, a sentence containing a simile, Two possible continuations, ending 0 and ending 1, where one reflects the correct figurative meaning and the other is often opposing or incorrect, and A binary label indicating the correct ending. This task evaluates whether language models can correctly interpret the intended figurative meaning of similes.
1.2 MAPS for Idiom
The MAPS dataset is designed to evaluate the understanding and interpretation of proverbs and idioms within conversational contexts. Proverbs and idioms are commonly used expressions that convey figurative meanings, which often differ significantly from their literal interpretations. Accurate comprehension requires models to infer the intended figurative meaning of these expressions based on context.
The MAPS dataset also presents a binary classification problem. Each instance, as illustrated in Table 2, consists of: proverb, a commonly used saying or idiomatic expression, conversation, a short dialogue where the proverb or idiom is used, two possible interpretations of the given conversation, one aligned with the figurative meaning of the proverb, and the other aligned with a literal or incorrect interpretation, and binary label (A or B) indicating the correct interpretation. This task assesses whether language models can accurately interpret the figurative meaning of idioms and proverbs when presented within a conversational context.
Appendix B Dataset Construction
@@ 2.1 Persian Simile Dataset
To create the Persian Simile dataset, we task GPT-4o with generating over 800 simile pairs, following the structure used in the MABL dataset. Each sample consists of a start phrase with two possible endings: one correctly reflecting the intended meaning and the other presenting an incorrect or opposite interpretation. Table 6 provides one example of used prompts.
| Your task is to generate pairs of sentences with opposite or very different meanings, both of which contain Persian similes. You can feel free to incorporate creativity into the similes, but also make sure that they’re something that could be understood by the speakers of the language that you are generating similes for, e.g., “this is as classic as pancakes for breakfast” to mean “this is classic” wouldn’t make sense for a culture in which pancakes aren’t traditionally eaten for breakfast. You can do this by thinking of a simile that conveys a certain meaning, and replacing the similative phrase with another similative phrase of the same type that conveys the opposite meaning. Here are some examples of similes to give you an idea of what we’re looking for: Please write the simile and two correct and incorrect meanings. 1. ت ه , . (Her hands are as strong and sturdy as the trunk of an oak tree.) ت ی / ت (Her hands are weak and frail / Her hands are strong and powerful) 2. ل ی (Her behavior is like a breeze blowing among the flowers) ./ . (Her behavior is aggressive and abrupt / His behavior is calm and composed) |